Measuring the Potential of a Nonideal Battery A battery with EMF 90.0 V has internal resistance Rb 39.10 2 Part A What is the reading VV of a voltmeter having total resistance Rv 3470 2 when it is placed across the terminals of the battery? Express your answer with three significant figures. Submit Hints My Answers Give Up Review Part Part B What is the maximum value that the ratio Rb /R may have if the percent error in the reading of the EMF of a battery is not to exceed 3.50% Express your answer with three significant figures. (Rb/Rv)max Submit Hints My Answers Give Up Review Part
Answer #1

Part A :The voltmeter reads the voltage which is equal to the voltage drop across the resistance Rv

Vv = I Rv   ------------------(1)

here "I" is the current in the circuit

Equivalent resistance of the circuit is

Req = Rb + Rv   ( Rb , Rv are in series )

Req = (9.10 \Omega + 470 \Omega ) = 479.10 \Omega .

Now the current in the circuit is I = EMF / (Req)

I = 90.0 V / ( 479.10 \Omega)

I = 0.1879 A.

now from (1) substituting all values we get

Vv = ( 0.1879 A ) ( 470 \Omega )

Vv = 88.31 V

Part B : Percent error of the reading of the EMF of a battery is

E = (V - Vv ) / V { E= Error , V =EMF = I ( Rb + Rv ) and Vv = I Rv }

E = [ I ( Rb + Rv ) -  I Rv ] / [ I ( Rb + Rv )]

E = [ Rb / (Rb + Rv ) ]

Rb E + Rv E = Rb

Rb (1-E) = E Rv

Rb / Rv = E / (1-E)

(Rb / Rv)max  = 0.036